OpenStudy (anonymous):

Let f(x)= –2x + 4 and g(x) = –6x – 7. Find f(x)– g(x). (2 points)

OpenStudy (ajprincess):

Find the difference between f(x) and g(x). f(x)-g(x)= –2x + 4-(–6x – 7). Does that help? @Acoustics95

OpenStudy (anonymous):

is that the solution? i dont understand @ajprincess

sam (.sam.):

That is not, simplify this –2x + 4-(–6x – 7)

OpenStudy (ajprincess):

Figured out the answer? @Acoustics95

OpenStudy (anonymous):

what would i do first to simplify that?

OpenStudy (ajprincess):

–2x + 4-(–6x – 7) =-2x+4+6x+7 what is -2x+6x=?

OpenStudy (anonymous):

4x?

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

4+7=?

OpenStudy (anonymous):

11

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

-2x+6x+7+4=?

OpenStudy (anonymous):

4x + 11?

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

that is ur answer:)

OpenStudy (anonymous):

Thank you very much @ajprincess

OpenStudy (anonymous):

:D

OpenStudy (ajprincess):

Welcome:)